By how much will GDP change, and will GDP be increasing, or will it be decreasing if: MPS = .1 and ↑ taxes = $11 billion
Answer #1

Answer

MPC=1-MPS=1-0.1=0.9

The tax multiplier=-MPC/(1-MPC)

=-0.9/(1-0.9)

=-9

The Change in GDP=change in taxes*tax multiplier

=11*(-9)

=-99 billion

the minus sign denotes the GDP decreases and it is decreased by $99 billion.